OpenStudy (anonymous):

amplitude:2 period:6pi phase shift:1/3 y=2sin(_x-_) how do i find the period for the equation?

OpenStudy (ash2326):

@ashleedean12 Suppose you have \[A\sin (\omega t+\theta)\] \[\omega=frequency\ in\ rad/s\] \[A=amplitude \] \[\theta= phase\ shift\] Relation between period T and \(\omega\) \[T=\frac{2 \pi}{\omega}\] here T= 6pi so \[6\pi=\frac{2\pi}{\omega}\] Find \(\omega\) from this ?

OpenStudy (anonymous):

would it be 3x?

OpenStudy (anonymous):

or 1/3?

OpenStudy (zpupster):

freq=2pi/period--rearrange per=2pi/frq

OpenStudy (ash2326):

w=1/3